Β A wire of length 0.050π‘š and carries a current 𝐼 is in a horizontal uniform magnetic field B of 2.50 𝑇 which is perpendicular to the rod. The weight of the wire is 0.80𝑁. What is the value of the current 𝐼 if the magnetic force on the wire equals its weight?



1
Expert's answer
2021-04-13T07:01:20-0400

To be given in question

Magnetic field=2.50 T

Force (F)=0.80N

Length (L)=0.050meter

"\theta=90°"

To be asked in question

Current "(I)" =?

We know that

"F=i(l\times B)"

"F=ilBsin\theta"

"sin90\u00b0=1"

"F=ilB"

Put value

"0.80=i\times0.050\times2.50"

"i=6.4 Ampear"